OpenStudy (anonymous):

You place a cup of 205oF coffee on a table in a room that is 72oF, and 10 minutes later, it is 195oF. Approximately how long will it be before the coffee is 180oF? Use Newton's law of cooling: \[T(t)=T _{A}+(T _{o}-T _{A})e ^{-kt}\] A. 15 minutes B. 25 minutes C. 1 hour D. 45 minutes

OpenStudy (anonymous):

Ta = 72 To=205

OpenStudy (anonymous):

use the same rule as before

OpenStudy (anonymous):

make sure you can use degrees F in this equation. i think u might have to convert to kelvin

OpenStudy (anonymous):

how would i do that?? :o

OpenStudy (anonymous):

i dont know if you have to do that or not, its been a long time since i have taken heat transfer

OpenStudy (anonymous):

hm, so just to clarify what I am doing, would i do ln(72/205) ?

OpenStudy (anonymous):

first you need to find k

OpenStudy (anonymous):

how do i do that?

OpenStudy (anonymous):

so you have 195 = 72 + (205-72)*e^(-k*10)

OpenStudy (anonymous):

(195-72)/(205-72) = e^(k*10)

OpenStudy (anonymous):

so take the ln of both sides, you get ln((195-72)/(205-72)) = k*10 solve for k

OpenStudy (anonymous):

then once you know k, u perform the same operation to find t for the second case, which is cooling down to 180 instead of 195